Can someone help me solve with this?
Solve the following.
sqrt(x+4)= x+4

Answer by rothauserc(4718)

sqrt(x+4)= x+4
square both sides of =
x+4 = x^2 +8x +16
x^2 +7x +12 = 0
factor
(x+4) * (x+3) = 0
x = -4 or -3